# Unifying KV Cache Compression for Large Language Models with LeanKV

### v2

@@ 3.2 Differentiated Token Importance
Tokens contribute to attention outputs with varying degrees of importance, as reflected by their attention scores. By exploiting these differences, we can apply finer-grained compression strategies that go beyond uniformly quantizing all tokens [70, 37] or exclusively pruning the least important ones [69, 10]. Specifically, we store progressively less important tokens at decreasing precision levels, ultimately leading to the pruning of the least important tokens.
To investigate the feasibility of this hierarchical approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones. Figure 2 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the general knowledge question answering benchmark MMLU [26] using the Llama3-8B model. For both pruning and the combined approach, approximately 50% of insignificant tokens are pruned. The resulting accuracy is nearly indistinguishable among the four approaches. However, the combination of 8-bit quantization and pruning achieves the lowest memory usage. The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
Figure 2: Accuracy and normalized KV cache memory usage of the Llama3-8B model on MMLU
3.3 Dynamic Sparsity Patterns
Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly [64, 69, 20, 35] or allocating it based on offline profiling [10]. However, LLMs exhibit dynamic sparsity patterns that vary across both attention heads and requests, which static memory allocation cannot adequately handle. To evaluate the sparsity of attention computation, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.
We begin by investigating dynamic sparsity patterns across layers. Using the first thousand articles from the Wikitext dataset, we evaluate the Llama3-8B model. Figure 3 illustrates the average number of critical tokens per layer, aggregated across all KV heads, required to retain 95% of the total attention score. Vertical bars represent the standard deviation, capturing variability across individual requests. Notably, the degree of sparsity varies considerably across layers, as indicated by differences in the number of critical tokens.
@@ 3.3 Dynamic Sparsity Patterns
These highly dynamic sparsity patterns, across layers, attention heads, and individual requests, underscore the necessity of adaptive memory management on a per-head and per-request basis.
@@ 3.4 Key Findings and Implications
Collectively, these insights guide the design of LeanKV, which integrates differentiated precision assignments for keys and values, hierarchical token importance handling, and per-request and per-head memory management.
@@ 4 KV Compression Policy
We present the KV compression policy of LeanKV that addresses the three levels of differentiation within the KV cache. First, to reflect the greater impact of keys on attention computation than values, we propose quantizing keys at a higher precision than values. For example, keys and values can be quantized to 8 and 4 bits (K8V4), or 4 and 2 bits (K4V2). Second, to account for the varying importance of tokens, we introduce a hierarchical compression strategy that classifies tokens into three significance levels: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at lower precision (e.g., K4V2), and the least important tokens are pruned. Finally, to address the dynamic sparsity patterns across requests and attention heads, we propose an adaptive memory management approach. Rather than imposing a fixed memory budget, our approach allows each attention head to determine its memory requirements dynamically based on its specific sparsity pattern.
@@
Prompt Phase. In the prompt phase, key and value vectors for all tokens in the prompt are computed. The compression policy then determines the appropriate precision for storing each token based on its significance. The significance of the $i^{th}$ token is calculated by averaging the $N-i$ attention scores it receives from subsequent tokens, where $N$ denotes the sequence length. In the case of GQA and MHA, scores from all attention heads associated with the KV head are aggregated using the maximum operation. To mitigate premature compression, the most recent $W$ tokens are always quantized at high precision, where $W$ is typically set to 64. For the remaining tokens, the policy determines the precision level of the $i^{th}$ token by comparing its significance score to the theoretical average $\frac{1}{i}$, based on the intuition that tokens with scores below average are less critical. Specifically, the $i^{th}$ token is quantized at high precision if its significance score exceeds $\frac{\alpha_{h}}{i}$, at low precision if its score lies within the interval $[\frac{\alpha_{l}}{i},\frac{\alpha_{h}}{i}]$, and is pruned otherwise. The parameters $\alpha_{l}$ and $\alpha_{h}$, which define the thresholds for high- and low-precision quantization, are determined offline through profiling on a calibration dataset. As a result of this hierarchical compression, the KV cache is conceptually divided into two parts: a high-precision section $KV_{h}$ and a low-precision section $KV_{l}$.
@@
Generation Phase. In the generation phase, a single token is compressed at each step, aligning with the autoregressive nature of the generation process. The most recent token is added to the recent window to prevent premature compression, while the earliest token $t_{c}$ in the window becomes a candidate for more aggressive compression. The compression procedure can be divided into two parts. First, token $t_{c}$ is quantized at either a high or low precision and added to the corresponding section of the KV cache, or it may be completely pruned. Next, if $t_{c}$ is quantized, the least significant token $t_{v}$ in the corresponding precision section of the KV cache is either re-quantized to lower precision or pruned, depending on its significance. Essentially, this policy establishes a smooth downgrading path for the less important token: rather than being pruned directly, it is first re-quantized to low precision, with pruning occurring only if it remains insignificant.
@@
The detailed procedure is outlined in Algorithm 1, following the same intuition and parameters as the prompt phase. Given the sequence length $N$, $t_{c}$ is quantized at high precision and added to $KV_{h}$, the high precision KV cache (line 6), if its significance exceeds $\frac{\alpha_{h}}{N}$. Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7). If $t_{v}$’s significance exceeds $\frac{\alpha_{h}}{N}$, it remains in $KV_{h}$; if it falls within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision KV cache (line 9); otherwise, $t_{v}$ is pruned. Similarly, if $t_{c}$’s significance lies within $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, it is quantized to low precision and added to $KV_{l}$ (line 14). The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $\frac{\alpha_{l}}{N}$ (line 17).
@@
Discussion. The proposed KV compression policy is highly extensible, allowing for the potential use of more than three precision levels. We adopt three precision levels primarily to minimize metadata overhead and improve system efficiency, as will be detailed in the following section. Additionally, we use a shared set of thresholds for all attention heads, as our empirical studies demonstrate that this approach is sufficient to capture the varying sparsity patterns across different heads. Notably, LeanKV can efficiently support more flexible compression policies, such as tuning thresholds for each head individually, which could potentially further optimize the balance between model accuracy and memory efficiency. We leave the exploration of such flexible policies for future work.
1:  Input: Parameters $\alpha_{h}$, $\alpha_{l}$; High & low precision $P_{h}$ $P_{l}$
2:  Input: Candidate token $t_{c}$; Sequence length $N$
3:  Input: High & low precision KV cache $KV_{h}$ $KV_{l}$
4:  Function: Significance $Score$; Quantization $Quant$;
@@
8:     if $T_{l}\leq$ Score($t_{v}$) $<T_{h}$ then
@@
10:     else if Score($t_{v}$) < $T_{l}$ then
@@
16:     if Score ($t_{v}$) < $T_{l}$ then
@@ 5 Memory Management
In this section, we describe LeanKV’s memory management mechanism which efficiently supports the three levels of differentiation in the KV cache. We first highlight the challenges posed by these differentiations in memory management, and then present parallel KV compaction, a novel memory management technique that effectively tackles the challenges.
@@ 5.1 Challenges
The Need for Flexible Paging. In PagedAttention [32], all tokens are stored at the same precision, allowing for a fixed page format. However, the differentiation of key and value vectors, along with varying token importance, introduces multiple precision levels both within individual tokens and across different tokens, making a fixed page format insufficient. Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage. Suppose memory is allocated to accommodate the highest precision K8V4, a token with K4V2 precision would waste 50% of the memory. Worse, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency.
Substantial KV Compaction Overhead. In PagedAttention, memory management complexity is $O(\#requests)$, as memory is partitioned uniformly across attention heads, with identical page IDs assigned to each head. In contrast, LeanKV must handle varying numbers of high- and low-precision tokens across different heads, mapping these irregular memory requirements to physical memory. We term this process KV compaction, whose complexity is $O(\#requests\times\#heads)$, significantly higher than that of PagedAttention. If KV compaction were performed sequentially on CPUs, its overhead could overshadow the performance gains from KV cache compression. Moreover, to track both high-precision and low-precision tokens, a straightforward approach would involve using separate data structures for each precision, resulting in increased metadata overhead.
@@ 5.2 Parallel KV Compaction
A detailed analysis of the KV compaction process reveals opportunities for parallelization and performance optimization. KV compaction can be divided into two phases: planning and coordination. In the planning phase, each head independently determines its memory allocation requirements. The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory. For instance, consider two attention heads require 16 and 32 bytes, respectively, as determined in the planning phase, with all GPU physical memory available. Then in the coordination phase, physical memory bytes 0 to 15 are allocated to the first head, and bytes 16 to 48 to the second. The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallel and well-suited to the GPU’s parallel compute capabilities. Parallelizing the coordination phase is more challenging, as it requires synchronization across heads. However, we observe that this phase can be parallelized effectively via parallel prefix sum [44], provided that the free memory region is contiguous.
@@
In parallel KV compaction, after each head determines the number of pages to be allocated or freed, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer, ensuring non-conflicting operations within the list. For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads. Similarly, for memory recycling, each head concurrently writes freed page IDs to its designated region, with the end pointer incremented by the total number of pages released.
Bidirectional Page Table. Similar to PagedAttention, LeanKV maintains a page table for each attention head, mapping each request to its corresponding list of page IDs. To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, LeanKV introduces a unified, GPU-resident data structure called the bidirectional page table. This structure efficiently supports the use of three precision levels for KV cache, as proposed in Section 4, minimizing the metadata overhead. In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload. The length of each page table entry is determined by the maximum sequence length divided by the tokens per high-precision page, ensuring no overflow since low-precision pages always contain more tokens than high-precision ones. This unified approach not only minimizes metadata overhead but also eliminates the need for separate lookups based on precision levels, thereby enhancing memory access efficiency during attention computations. The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama-2-7B, which has 32 layers and 32 attention heads per layer, the total size of all bidirectional page tables is only 64 MB. By contrast, the KV cache for a single request occupies 2 GB.
@@ 5.3 KV Compaction Workflow
We delve into the KV compaction workflow during the prompt and generation phases, illustrating how the three data structures interact to enable differentiated KV cache compression. Notably, KV compaction is executed once per inference step for all requests in the batch and all layers in the LLM. This design ensures sufficient parallelism for efficient GPU execution of parallel KV compaction, while amortizing the associated GPU kernel launch overheads.
@@
Prompt Phase. Figure 5 illustrates the KV compaction workflow during the prompt phase using an example request with eight tokens. In this setting, one high-precision page can store two tokens, and one low-precision page can store four. Since the exact number of high- and low-precision pages required by each head is unknown prior to running the compression algorithm, we conservatively allocate four unified pages per head, assuming that all tokens will be stored at high precision. Each head then independently applies the KV cache compression algorithm to determine which tokens to store at high-precision, low-precision, or to prune. This step constitutes the planning phase, during which each head calculates its specific memory needs based on its unique sparsity pattern. In this example, the first head uses one high-precision page, while the second head uses two; both heads also require one low-precision page. High-precision pages are allocated from the left side of the bidirectional page table, while low-precision pages are allocated from the right side. Following the planning phase, any unused pages are marked for recycling, triggering the coordination phase. Here, we employ a parallel prefix-sum-based approach, as described in the previous section, to facilitate efficient and conflict-free page recycling across heads.

### v3

@@ 3.2 Differentiated Token Importance
Tokens contribute to attention outputs with varying degrees of importance, as reflected by their attention scores. By exploiting these differences, we can apply finer-grained compression strategies that go beyond uniformly quantizing all tokens [77, 44] or exclusively pruning the least important ones [76, 11].
Figure 3 shows the per-token attention scores in the $8^{th}$ layer, selected as a representative layer of Llama3-8B [20], on a sequence randomly sampled from the Wikitext dataset [49]. Similar sparsity patterns are observed across the other layers as well. This non-uniform distribution of token importance motivates a hierarchical compression strategy that allocates memory based on token significance: high-precision storage (e.g., K8V4) for the most important tokens, lower precision (e.g., K4V2) for moderately important tokens, and pruning for the least important. This hierarchical approach enables a smoother trade-off between memory savings and generation quality than uniform quantization methods (e.g., Atom [77], QServe [44]) and pruning-only approaches (e.g., H2O [76], SnapKV [42]).
Figure 3: Per-token attention scores in the $8^{th}$ layer of Llama3-8B on a sequence randomly sampled from Wikitext.
3.3 Per-head Dynamic Sparsity Patterns
LLMs exhibit dynamic sparsity patterns that vary across both attention heads and requests: the number of critical tokens can differ not only between heads but also for the same head under different requests. To characterize the sparsity of attention, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.
We begin by investigating dynamic sparsity patterns across layers, and evaluate the Llama3-8B model with the first thousand articles from the Wikitext dataset. Figure 4 illustrates the average number of critical tokens per layer, aggregated across all KV heads, required to retain 95% of the total attention score. Vertical bars represent the standard deviation, capturing variability across individual requests. Notably, the degree of sparsity varies considerably across layers, as indicated by differences in the number of critical tokens.
@@ 3.4 Main Insights and Implications
Collectively, these highly dynamic sparsity patterns, across layers, attention heads, and individual requests, underscore the necessity of adaptive memory management on a per-head and per-request basis. They motivate the design of DiffKV, which integrates differentiated precisions for keys and values, hierarchical compression based on importance, and per-request and per-head memory management.
Contributions of DiffKV. DiffKV expands the design space for KV cache compression by jointly exploiting the three levels of differentiation in KV cache, enabling higher compression ratios with minimal quality degradation. Compared to prior work that explores importance-based KV cache pruning [76, 42, 5, 65] and mixed-precision quantization [77, 19, 30], DiffKV introduces the following key innovations:
  • Request-aware differentiated KV quantization. DiffKV adaptively adjusts the mix of high- and low-precision tokens (e.g., K8V4, K4V2) on a per-request basis (Section 4), whereas prior work [77, 19, 5] typically applies a single static configuration across all requests.
  • Sequence length-adaptive importance estimation. Diff-KV modulates the proportion of important tokens based on sequence length, retaining more tokens in shorter sequences to preserve quality, while enabling more aggressive compression for longer ones (Section 4). Prior approaches [76, 42, 5] typically use a fixed token retention ratio, regardless of sequence length.
  • Per-head dynamic sparsity exploitation. DiffKV dynamically adjusts memory allocation per head, based on observed sparsity patterns (Section 3.3 & 4), whereas existing methods allocate memory uniformly across attention heads, missing optimization opportunities.
  • Scalable on-GPU memory management. DiffKV introduces an efficient on-GPU memory manager that handles irregular per-head memory allocation patterns, effectively translating memory savings into performance gains (Section 5).
@@ 4 KV Compression Policy
We present the KV compression policy of DiffKV, designed to address the three levels of differentiation within the KV cache. First, to reflect the greater impact of keys on attention computation compared to values, we propose quantizing keys at a higher precision than values. For example, keys and values can be quantized to 8 and 4 bits (K8V4), or 4 and 2 bits (K4V2). Second, to account for the varying importance of tokens, we introduce a hierarchical compression strategy that classifies tokens into three significance levels: the most important tokens are quantized at high precision (e.g., K8V4), moderately important tokens at lower precision (e.g., K4V2), and the least important tokens are pruned entirely. This strategy is both request-aware and sequence-length adaptive. It dynamically adjusts the mix of high- and low-precision tokens per request to reflect their varying sparsity patterns. Moreover, it retains a larger fraction of tokens at high precision for short sequences to preserve quality, while applying low precision quantization and pruning more aggressively for long sequences to maximize memory savings.
Finally, to address the dynamic attention sparsity patterns across requests and attention heads, we propose an adaptive memory management approach. Rather than imposing a fixed memory budget, our approach allows each attention head to determine its memory requirements dynamically based on its specific sparsity pattern.
@@
Prompt Phase. In the prompt phase, key and value vectors for all tokens in the prompt are computed. The compression policy then determines the appropriate precision for storing each token based on its significance. The significance of the $i^{th}$ token is calculated by averaging the NiN-i attention scores it receives from subsequent tokens, where $N$ denotes the prompt sequence length. In the case of GQA and MHA, scores from all attention heads associated with the KV head are aggregated using the maximum operation. To mitigate premature compression, the most recent $W$ tokens are always quantized at high precision, where $W$ is typically set to 64. For the remaining tokens, the policy determines the precision level of the $i^{th}$ token in a sequence-length adaptive manner, by comparing its significance score to the theoretical average $\frac{1}{i}$, based on the intuition that tokens with scores below average are less critical. Specifically, the $i^{th}$ token is quantized at high precision if its significance score exceeds $\frac{\alpha_{h}}{i}$, at low precision if its score lies within the interval $\frac{\alpha_{l}}{i},\frac{\alpha_{h}}{i}]$, and is pruned otherwise. The parameters $\alpha_{l}$ and $\alpha_{h}$, which define the thresholds for high- and low-precision quantization, are determined offline through profiling on a calibration dataset. As a result of this hierarchical compression, the KV cache is conceptually divided into two parts: a high-precision section $KV_{h}$ and a low-precision section $KV_{l}$.
@@
Generation Phase. In the generation phase, a single token is compressed at each step, aligning with the autoregressive nature of the generation process. The most recent token is added to the recent window to prevent premature compression, while the earliest token $t_{c}$ in the window becomes a candidate for more aggressive compression. The compression procedure can be divided into two parts. First, token $t_{c}$ is quantized at either a high or low precision and added to the corresponding section of the KV cache, or it may be completely pruned. Next, if $t_{c}$ is quantized, the least significant token $t_{v}$ in the corresponding precision section of the KV cache is considered for further downgrading: it may be re-quantized to lower precision or pruned, depending on its significance. Essentially, this policy establishes a smooth downgrading path for the less important token: rather than being pruned directly, it is first re-quantized to low precision, with pruning occurring only if it remains insignificant.
@@
The detailed procedure is outlined in Algorithm 1, following the same intuition and parameters as the prompt phase. Given the sequence length $N$, $t_{c}$ is quantized at high precision and added to $KV_{h}$, the high precision KV cache (line 6), if its significance exceeds $\frac{\alpha_{h}}{N}$. Subsequently, the least significant token in $KV_{h}$ is identified as the victim token $t_{v}$ for more aggressive compression (line 7). If $t_{v}$’s significance exceeds $\frac{\alpha_{h}}{N}$, it remains in $KV_{h}$; if it falls within $\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, $t_{v}$ is re-quantized to low precision and moved to $KV_{l}$, the low precision KV cache (line 9); otherwise, $t_{v}$ is pruned. Similarly, if $t_{c}$’s significance lies within $\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, it is quantized to low precision and added to $KV_{l}$ (line 14). The least important token in $KV_{l}$ is designated as the victim $t_{v}$ (line 15), and is further pruned if its significance falls below $\frac{\alpha_{l}}{N}$ (line 17).
@@
Discussion. The proposed KV compression policy is highly extensible, allowing for the potential use of more than two quantization precision levels. We adopt two quantization precision levels primarily to minimize metadata overhead and improve system efficiency, as will be detailed in the following section. In Section 7.2, we empirically show that the two-level scheme (K8V4-K4V2) achieves near-lossless generation quality across multiple models and benchmarks. We use a shared set of thresholds for all attention heads, as our empirical studies demonstrate that this approach is sufficient to capture the varying sparsity patterns across different heads. Notably, DiffKV can efficiently support more flexible compression policies, such as tuning thresholds for each head individually, which could potentially further optimize the balance between model accuracy and memory efficiency. We leave the exploration of such flexible policies for future work.
1:Input: Parameters $\alpha_{h}$, $\alpha_{l}$; High & low precision $P_{h}$ & $P_{l}$
2:Input: Candidate token $t_{c}$; Sequence length NN
3:Input: High & low precision KV cache $KV_{h}$ & $KV_{l}$
4:Function: Significance ScoreScore; Quantization QuantQuant;
@@
8:  if $\frac{\alpha_{l}}{N}\leq$ Score($t_{v}$) < $\frac{\alpha_{h}}{N}$ then
@@
10:  else if Score($t_{v}$) < $\frac{\alpha_{l}}{N}$ then
@@
16:  if Score ($t_{v}$) < $\frac{\alpha_{l}}{N}$ then
@@ 5 Memory Management
In this section, we describe DiffKV’s memory management mechanism which efficiently supports the three levels of differentiation in the KV cache. We first highlight the challenges posed by these differentiations in memory management, and then present parallel KV compaction, a novel memory management technique that effectively tackles the challenges.
@@ 5.1 Challenges
Flexible Paging. In PagedAttention [39], all tokens are stored at the same precision, allowing for a fixed page format. However, the differentiation of key and value vectors, along with varying token importance, introduces multiple precision levels both within and across tokens, making a fixed page format insufficient. Specifically, a fixed page format requires conservatively allocating high-precision slots for all tokens regardless of their actual precisions, leading to considerable memory wastage. Suppose memory is allocated to accommodate the highest precision K8V4, a token with K4V2 precision would waste 50% of the memory. Worse, such a fixed page format results in misaligned memory accesses, which hinder memory bandwidth utilization and reduce overall computational efficiency.
KV Compaction Scalability. At each LLM inference step, the memory management complexity is O(#requests)O(\#requests) for PagedAttention, as memory is partitioned uniformly among attention heads. In contrast, DiffKV must handle varying numbers of high- and low-precision tokens across different heads, mapping these heterogeneous memory requirements to physical memory. We term this process KV compaction, whose complexity is $O(\#requests\times\#heads)$ per inference step. This poses significant scalability challenges for memory management. Given that model execution time per step is on the order of tens of milliseconds, improper memory management could render the overhead of dynamic allocation prohibitive, negating the benefits of KV cache compression. Moreover, managing both high-precision and low-precision tokens would require separate data structures for each precision, resulting in increased metadata overhead.
@@ 5.2 Parallel KV Compaction
A detailed analysis of the KV compaction process reveals opportunities for parallelization. KV compaction can be divided into two phases: planning and coordination. In the planning phase, each attention head independently determines its memory allocation requirements. The subsequent coordination phase synchronizes these per-head requirements and maps them to the GPU’s physical memory. The primary source of increased memory management complexity lies in the planning phase, which is perfectly parallelizable and well-suited to the GPU’s parallel compute capabilities. Parallelizing the coordination phase is more challenging, as it requires synchronization across heads. However, we observe that this phase can be parallelized effectively via parallel prefix sum [51], provided that the free memory region is contiguous.
@@
In parallel KV compaction, after each head determines the number of pages to be allocated or freed, a parallel prefix sum operation computes a unique offset for each head relative to the start or end pointer. This step converts per-head memory demands into a globally consistent schedule, where each head is assigned a disjoint region of the free page list to read from (for allocation) or write to (for recycling), ensuring non-conflicting operations within the list. For memory allocation, each head concurrently retrieves its new page IDs from its designated region in the list, with the start pointer incremented by the cumulative number of pages required across all heads. Similarly, for memory recycling, each head writes freed page IDs to its designated region, with the end pointer incremented by the total number of pages released.
Bidirectional Page Table. Similar to PagedAttention, Diff-KV maintains a page table for each attention head, mapping each request to its corresponding list of page IDs. To avoid the doubled metadata overhead associated with maintaining separate page tables for high- and low-precision pages, DiffKV introduces a unified, GPU-resident data structure called the bidirectional page table. This structure efficiently supports the use of two quantization precision levels for KV cache, as proposed in Section 4, minimizing the metadata overhead. In each entry of the bidirectional page table, high-precision page IDs grow from the left side of the list, while low-precision page IDs grow from the right, dynamically adapting to the precision requirements of the workload. The length of each page table entry is determined by the maximum sequence length divided by the tokens per high-precision page, ensuring no overflow since low-precision pages always contain more tokens than high-precision ones. This unified approach not only minimizes metadata overhead but also eliminates the need for separate lookups based on precision levels, thereby enhancing memory access efficiency during attention computations. The memory overhead of the bidirectional page table is minimal: for example, with a batch size of 128 on Llama3-8B, which has 32 layers and 8 KV heads per layer, the total size of all bidirectional page tables is only 32 MB. By contrast, the KV cache for a single request occupies 1 GB.
@@ 5.3 KV Compaction Workflow
We delve into the KV compaction workflow during the prompt and generation phases, illustrating how the three data structures interact to enable differentiated KV cache compression. Notably, KV compaction is executed once per inference step for all requests in the batch and all layers in the LLM. This design ensures sufficient parallelism for efficient GPU execution, while amortizing the associated GPU kernel launch overheads.
@@
Prompt Phase. Figure 6 illustrates the KV compaction workflow during the prompt phase for an example request with eight tokens. In this example, one high-precision page stores two tokens, and one low-precision page stores four. Since the exact numbers of high- and low-precision pages required by each head are unknown a priori, we conservatively allocate four unified pages per head, assuming all tokens are stored at high precision. Pages 5-8 and 9-12 are thus assigned to Head A and B, respectively, as shown in their page tables. The end pointer of circular free page list advances to Page 13, marking Pages 5-12 as allocated.
Next comes the planning phase, during which each head independently applies the KV cache compression algorithm and calculates its specific memory requirements. In this example, Head A uses one high-precision page, while Head B uses two; both heads also require one low-precision page. We allocate high-precision pages left-to-right in the bidirectional page table and low-precision pages right-to-left: Head A uses Page 5 (high-precision) and Page 8 (low-precision), while Head B uses Pages 9–10 (high-precision) and Page 12 (low-precision).
Following the planning phase, the coordination phase finally reclaims unused pages (Pages 6-7 from Head A and Page 11 from Head B) via a parallel prefix-sum. These recycled pages are appended to the circular free page list; since the end pointer already points to the list’s tail, it wraps around to the head to accommodate the recycled pages.
@@
Supporting Additional Precision Levels. The proposed memory management system naturally extends to support more than two precision levels, by incorporating additional page tables. For example, accommodating three precision levels can be achieved by combining a unidirectional page table with a bidirectional one. Similarly, four precision levels can be supported using two bidirectional page tables. This composable design preserves memory efficiency while providing the flexibility needed to handle increasingly fine-grained precision differentiation.